# Using Covid-19 Response Policy to Estimate Open Water Swim Drafting Effects in Triathlon

### v1

@@ 5. Compute individual drafting positions within each cluster (swim group):
Findings II.
Drafting is beneficial compared to the leader only when an athlete is in at least the third drafting position within a swim group constructed based on clustered swim-out times. Drafting positions 1 and 2 are therefore strategic primarily as positions to minimize fatigue and strain before the subsequent bike leg in triathlon races. Since swim groups are constructed based on swim-out times, there may be some degree of endogeneity in the estimates. This could arise because athletes who manage to catch up to a swim group toward the end of the swim segment might influence their position within the swim group, potentially leading to biased results. To mitigate some of this endogeneity, athlete and event fixed effects are employed.
@@ Conclusion
This study provides the first large-scale causal estimate of open-water swim drafting effects in real-world triathlon race settings. Leveraging a natural experiment induced by staggered race starts during the COVID-19 pandemic, we find that drafting benefits were statistically insignificant in 2020 but re-emerged post-pandemic, albeit at slightly reduced levels. Our results indicate that drafting only becomes beneficial from the third trailing position onward, while earlier positions primarily serve to minimize fatigue or strategic purposes. By employing athlete and event fixed effects, I address potential heterogeneity concerns, and the observed inverse decay of drafting benefits helps to partially mitigate concerns related to endogeneity, reverse causality and omitted variable bias. These findings offer valuable insights into optimal swim strategies for endurance athletes and race organizers.
Disclosure Statement
The authors declare no conflicts of interest. No funding for this research was provided. The study adheres to ethical guidelines, and no human subjects were directly involved in the research.
@@
Using Covid-19 Response Policy to Estimate Open Water Swim Drafting Effects in Triathlon / The Causal Effect of Higher-Order Water Swim Drafting Positions: Evidence from COVID-19 Response Policies

### v2

@@ Results
The Effect of Higher-Order Drafting Positions
Drafting is beneficial compared to the leader (based purely on swim-out segment performance time) only when an athlete is in at least the third drafting position within a swim group constructed based on clustered swim-out times. Drafting positions 1 and 2 are therefore strategic primarily as positions to minimize fatigue and strain before the subsequent bike leg in triathlon races. Delextrat et al. (2003) Since swim groups are constructed based on swim-out times, there may be some degree of endogeneity in the estimates. This could arise because athletes who manage to catch up to a swim group toward the end of the swim segment might influence their position within the swim group, potentially leading to biased results. To mitigate some of this endogeneity, athlete and event fixed effects are employed.
@@
Discussion
While I provide the first causal estimates of open-water swim drafting effects, several limitations should be acknowledged.
First, the estimation approach relies on inferred swim group structures based on clustered Swim-Out times, which may introduce measurement error. The classification of drafters and leaders remains an approximation rather than a direct observation.
Second, the staggered race starts during the COVID-19 pandemic serve as a natural experiment. However, unobserved factors, such as changes in athlete fitness levels, race-day conditions, or competition field strength, may have influenced drafting benefits across different periods, potentially confounding causal interpretations. This might have led to the only weak significance (at the 10% level) of the interaction term Pre-period $\times$ Drafter in Table 3.
Third, reverse causality remains a concern despite the inclusion of athlete- and event-fixed effects. More capable athletes may self-select into different drafting positions, catch up to swim groups over the course of a race, or utilize higher passive drag depending on their fitness level on race day. These dynamics could result in endogenous group formations, potentially biasing the estimated effects.
Finally, course-specific factors such as buoy placements, water currents, and water temperature—which influence wetsuit usage regulations—are not explicitly controlled for in the models. These factors may affect drafting effectiveness but are difficult to quantify systematically across different race events.
@@ Conclusion
This study provides the first large-scale causal estimate of open-water swim drafting effects in real-world triathlon race settings. Leveraging a natural experiment induced by staggered race starts during the COVID-19 pandemic, I find that drafting benefits were statistically insignificant in 2020 but re-emerged post-pandemic, albeit at slightly reduced levels. Our results indicate that drafting only becomes beneficial from the third trailing position onward, while earlier positions primarily serve to minimize fatigue or strategic purposes. By employing athlete and event fixed effects, I address potential heterogeneity concerns, and the observed inverse decay of drafting benefits helps to partially mitigate concerns related to endogeneity, reverse causality,, and omitted variable bias. These findings offer valuable insights into optimal swim strategies for endurance athletes and race organizers.
Declaration of Interest
The authors declare no conflicts of interest.
Funding Information
No funding for this research was provided.
Confirmation of Ethical Compliance
The study adheres to ethical guidelines, and no human subjects were directly involved in the research.
CRediT Author Statement
Conceptualization: Felix Reichel;
Methodology: Felix Reichel;
Formal analysis: Felix Reichel;
Investigation: Felix Reichel;
Data curation: Manuel Poeter MSc, Mag. Thomas Doblhammer;
Writing – original draft: Felix Reichel;
Writing – review & editing: Felix Reichel;
Visualization: Felix Reichel;
Supervision: Felix Reichel.
@@
Using Covid-19 Response Policy to Estimate Open Water Swim Drafting Effects in Triathlon